# LLAMAFUZZ: Large Language Model Enhanced Greybox Fuzzing

### v3

@@ C.2 Coverage-guided greybox fuzzing
As the fuzzing loop continues, more code branches will be reached, which holds the potential to trigger a bug Wang et al. (2019a).
C.3 Fuzzing for structured data
@@
Fuzzers for structured data Wang et al. (2019b); Le et al. (2021); Mallissery and Wu (2023); Zhang et al. (2024); Meng et al. (2024); Koffi et al. (2024); Shi et al. (2023) can accurately identify the target input format. They can generate test cases that maintain the consistency of the format. This approach ensures that the generated test cases are not only valid but also effective in triggering and exploring potential vulnerabilities or issues within the application. Three grammar-aware mutation operators have been found to be particularly effective in uncovering deep bugs Srivastava and Payer (2021); Aschermann et al. (2019): random mutation, which involves selecting a random non-leaf non-terminal node and creating a new context-free grammars derivation subtree. Random recursive unrolling, which finds recursive production rules and expands them up to $n$ times. Splicing, which combines two inputs while preserving their syntactic validity. In addition, Langfuzz Holler et al. (2012) combines grammar-based fuzz testing and reusing project-specific issue-related fragments, maintaining the integrity of format and having a higher chance to cause new problems than random input. QuickFuzz Grieco et al. (2016) leverages Haskell’s QuickCheck and the Hackage to fuzz structured data. This integration, combined with conventional bit-level mutational fuzzers, negates the need for an external set of input files and eliminates the requirement to develop specific models for the file types being tested. Alternatively, WEIZZ Fioraldi et al. (2020a) employs a chunk-based mutator to generate and mutate inputs for unknown binary formats. Nevertheless, WEIZZ struggles to handle grammar-based formats such as JSON, XML, and programming languages.
C.4 Augment fuzzing through machine learning
@@
C. Pérez Pérez et al. (2024) explored the ability of Compressed-Language Models (CLMs) to interpret files compressed by standard file formats. Their findings revealed that CLMs are capable of understanding the semantics of compressed data directly from the byte streams, opening a new path for processing raw compressed files. In a related study, CHATFUZZHu et al. (2023) investigates the mutation capabilities of LLM on text-based seeds, achieving 12.77% edge coverage improvement over the SOTA greybox fuzzer (AFL++). Similarly, SmartSeed Lyu et al. (2018) combines deep learning models to generate new inputs for evaluating 12 different applications.
@@
Prior work Eom et al. (2024) integrates an LLM-based mutator with a reinforcement learning approach, utilizing the Term Frequency-Inverse Document Frequency technique to develop a weighted coverage map. This method capitalizes on coverage feedback to enhance the effectiveness of the mutation process. Similarly, Xia et al. Xia et al. (2024) introduce an auto-prompting phase that employs LLMs to produce and mutate test cases across six programming languages. Their findings indicate that LLMs can surpass the coverage achieved by cutting-edge tools.
@@
Additionally, WhiteFox Yang et al. (2023) employs dual LLMs within their framework: one analyzes low-level optimization source code to inform optimization strategies, while the other generates test programs based on this analysis. CHATAFL Meng et al. (2024) utilizes LLMs to understand protocol message types and assesses their ability to identify "states" in stateful protocol implementations. LLM4FUZZ Shou et al. (2024) leverages LLMs to guide fuzzers towards more critical code areas and input sequences that are more likely to reveal vulnerabilities, showcasing the potential of LLMs in prioritizing and refining fuzzing efforts.
C.5 Large Language Model
In recent studies, pre-trained Large Language Models (LLMs) have shown impressive performance on natural language tasks, including Natural language understanding, reasoning, natural language generation, multilingual, and factuality Chang et al. (2023); Menon and Vondrick (2022); Madani et al. (2023); Li et al. (2022).
Utilizing unsupervised learning, Large Language Models are pre-trained on extensive textual data, enabling LLM with a broad range of knowledge. Additionally, with billions or trillions of parameters, LLM can not only capture the patterns in context but also understand the textual data at a deeper level such as format and chunk information within files. Such capabilities have facilitated LLMs to exhibit remarkable competencies beyond traditional Natural Language Processing tasks. Evidence of their versatility includes visual classification Menon and Vondrick (2022), protein sequence generation Madani et al. (2023), code generation Li et al. (2022).
Building upon this versatile foundation. This inherent capability to interpret and process different data structures renders LLMs particularly effective in the mutation stage of fuzzing processes. CHATFUZZ Hu et al. (2023) employs LLMs to directly generate seeds, though its application is limited to text-based target programs such as JSON and XML. Moreover, Pérez et al. demonstrate that Compressed-Language Models can understand files compressed by Compressed File Formats.
Appendix D Data Availability
D.1 LLAMAFUZZ
The LLAMAFUZZ system comprises two primary components: a greybox fuzzer and a large language model (LLM). The greybox fuzzer (AFL++) in LLAMAFUZZ is available in two versions to cater to different types of seeds:
Binary Seeds Version: This version of the fuzzer is designed to handle binary seeds, providing efficient and effective fuzzing for binary-based targets. https://anonymous.4open.science/r/AFLplusplus-binary-seeds
Text-Based Seeds Version: This version is tailored for text-based seeds, ensuring comprehensive coverage and mutation for text-based targets. https://anonymous.4open.science/r/AFLplusplus-text-based
The link below provided the LLM component of LLAMAFUZZ:
D.2 Experiment benchmarks
We have extended the base version of benchmark frameworks to support LLAMAFUZZ. This includes added functionality to seamlessly integrate the greybox fuzzer and LLM into the benchmarking process.

### v4

@@ 6.2. Coverage-guided greybox fuzzing
As the fuzzing loop continues, more code branches will be reached, which holds the potential to trigger a bug (Wang et al., 2019a).
6.3. Fuzzing for structured data
@@
Fuzzers for structured data (Wang et al., 2019b; Le et al., 2021; Mallissery and Wu, 2023; Zhang et al., 2024; Meng et al., 2024; Koffi et al., 2024; Shi et al., 2023) can accurately identify the target input format. They can generate test cases that maintain the consistency of the format. This approach ensures that the generated test cases are not only valid but also effective in triggering and exploring potential vulnerabilities or issues within the application. Three grammar-aware mutation operators have been found to be particularly effective in uncovering deep bugs (Srivastava and Payer, 2021; Aschermann et al., 2019): random mutation, which involves selecting a random non-leaf non-terminal node and creating a new context-free grammars derivation subtree. Random recursive unrolling, which finds recursive production rules and expands them up to $n$ times. Splicing, which combines two inputs while preserving their syntactic validity. In addition, Langfuzz (Holler et al., 2012) combines grammar-based fuzz testing and reusing project-specific issue-related fragments, maintaining the integrity of format and having a higher chance to cause new problems than random input. QuickFuzz (Grieco et al., 2016) leverages Haskell’s QuickCheck and the Hackage to fuzz structured data. This integration, combined with conventional bit-level mutational fuzzers, negates the need for an external set of input files and eliminates the requirement to develop specific models for the file types being tested. Alternatively, WEIZZ (Fioraldi et al., 2020a) employs a chunk-based mutator to generate and mutate inputs for unknown binary formats. Nevertheless, WEIZZ struggles to handle grammar-based formats such as JSON, XML, and programming languages.
6.4. Augment fuzzing through machine learning
@@
C. Pérez (Pérez et al., 2024) explored the ability of Compressed-Language Models (CLMs) to interpret files compressed by standard file formats. Their findings revealed that CLMs are capable of understanding the semantics of compressed data directly from the byte streams, opening a new path for processing raw compressed files. In a related study, CHATFUZZ(Hu et al., 2023) investigates the mutation capabilities of LLM on text-based seeds, achieving 12.77% edge coverage improvement over the SOTA greybox fuzzer (AFL++). Similarly, SmartSeed (Lyu et al., 2018) combines deep learning models to generate new inputs for evaluating 12 different applications.
@@
Prior work (Eom et al., 2024) integrates an LLM-based mutator with a reinforcement learning approach, utilizing the Term Frequency-Inverse Document Frequency technique to develop a weighted coverage map. This method capitalizes on coverage feedback to enhance the effectiveness of the mutation process. Similarly, Xia et al. (Xia et al., 2024) introduce an auto-prompting phase that employs LLMs to produce and mutate test cases across six programming languages. Their findings indicate that LLMs can surpass the coverage achieved by cutting-edge tools.
@@
Additionally, WhiteFox (Yang et al., 2023) employs dual LLMs within their framework: one analyzes low-level optimization source code to inform optimization strategies, while the other generates test programs based on this analysis. CHATAFL (Meng et al., 2024) utilizes LLMs to understand protocol message types and assesses their ability to identify ”states” in stateful protocol implementations. LLM4FUZZ (Shou et al., 2024) leverages LLMs to guide fuzzers towards more critical code areas and input sequences that are more likely to reveal vulnerabilities, showcasing the potential of LLMs in prioritizing and refining fuzzing efforts.
7. Conclusion